OpenStudy (anonymous):

A colony of insects grows exponentially from 100 to 146 in one week. How many DAYS does it take for the population to reach 1300?

OpenStudy (amistre64):

146/100 = 1.46; so it inceases at 146% in 7 days eh....

OpenStudy (amistre64):

or is that 46% :)

OpenStudy (yuki):

You do something similar with the ozone one let me just leave the formula since amistre can help you with the concept\[N_t = N_0(1.46)^{t/7}\]

OpenStudy (amistre64):

7 days = 1/2x{0} + x{0} = x{n+1} x{1} = 100(1.5) = 150 x{2} = 150(1.5) = ......is that what i was missing last time the 1?

OpenStudy (amistre64):

...... good job yuki :)

OpenStudy (anonymous):

Ok, so the equation should look like 1300= 100(1.46)^(t/7).....right?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

moving on, I should get to (log 13/log 1.46) = t/7.......right?

OpenStudy (anonymous):

Which gives me 47.44, so the answer is c?

OpenStudy (anonymous):

c is 47, sorry I forgot its not on here

OpenStudy (yuki):

remidia, since I don't know what the answer choices are I cannot say this by 100%, but it takes exactly 7*ln13/ln1.46 days to reach 1300. Which means that if it is any less than that you still haven't reached 1300. so, if there, by any chance, were two solutions 47 days or 48 days, make sure to pick 48 days, ok ?

OpenStudy (anonymous):

Ok, thank you!